OpenStudy (anonymous):

Solve for y. 2y – 16 = –4 A. –10 B. –18 C. 6 D. 14

OpenStudy (anonymous):

@Destinymasha

OpenStudy (destinymasha):

substitute....

OpenStudy (anonymous):

its me destiny

OpenStudy (anonymous):

trenton,maurice

OpenStudy (anonymous):

dont remember???

OpenStudy (destinymasha):

yup...

OpenStudy (destinymasha):

answer is C

OpenStudy (destinymasha):

HI LOVE!!!

OpenStudy (anonymous):

you know your math too?

OpenStudy (anonymous):

your too smart

OpenStudy (anonymous):

sorry bout yesterday had to clean up my room lol

OpenStudy (destinymasha):

2 [6] = 12, 12 - 16 = -4 -im horrible at math, hate it ...lol

OpenStudy (destinymasha):

omg, can you message me?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Solve for x. 8 – 3x – x = 12 A. –2 B. –5 C. –1 D. –10

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

but answer this first missy

OpenStudy (destinymasha):

ok

OpenStudy (anonymous):

:Solve for x. 8 – 3x – x = 12 A. –2 B. –5 C. –1 D. –10

OpenStudy (anonymous):

brb Breakfeast! :DDDDDDDDDDDDDDDdd

OpenStudy (destinymasha):

C [ -1 ]

OpenStudy (anonymous):

@Destinymasha

OpenStudy (destinymasha):

back?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

i hate BACON

OpenStudy (destinymasha):

well... there's the answer ^ ^

OpenStudy (destinymasha):

im sorry, lol - me too

OpenStudy (anonymous):

:Identify the solution set of the inequality using the given replacement set. x < –7; {–10, –7.2, –7, –6.9, 6.5, 9} A. {–7, –6.9, 6.5} B. {–6.9, 6.5} C. {–10, –7.2} D. {–10, –7.2, –7}

OpenStudy (anonymous):

wow really thats crazy

OpenStudy (destinymasha):

it means x has to be less than or equal to 7 ...

OpenStudy (destinymasha):

so....C. {–10, –7.2}

OpenStudy (destinymasha):

those are both less than -7

OpenStudy (destinymasha):

SORRY - less than -7 ...

OpenStudy (anonymous):

so um what is tthe answer to this one?

OpenStudy (anonymous):

@Destinymasha !!!

OpenStudy (anonymous):

@Destinymasha

OpenStudy (destinymasha):

sorry love, my stupid computer....

OpenStudy (anonymous):

@Destinymasha

OpenStudy (anonymous):

your only suppose to help me duh